A spherical balloon is 40 ft in diameter and surrounded by air at 60 0F and

29.92 in Hgabs. If the balloon is filled with hydrogen at a temperature of 70

0Fand atmospheric pressure, what total load can it lift? R of hydrogen is

766.54 ft-lbf/lbm-R. (2381 lb)


1
Expert's answer
2022-02-15T10:59:30-0500

We know that

"T=60F=60+460=520R"

"d=40ft"

"V=\frac{4}{3}\times\pi r^3=\frac{4}{3}\times3.14\times20^3=33510.32ft^3"

"P=29.92 inHg abs"

"PV=mRT"

"m_{air}=\frac{PV}{RT}"


"m_{air}=\frac{29.92\times33510.32}{766.54\times520}=2557.4201lb"

Pul value solve it


"m_{air}=2557.4201lb"

"M_{H}=174.6014lb"

Now "M_{lift}=m_{air}-M_{H}"


"M_{lift}=2557.4201-174.6014=2382.8187lb"